OpenStudy (anonymous):

how do i find the maximum of y=.75x^2 - .04x^3

OpenStudy (anonymous):

Maximum of a function occurs when dy/dx=0 dy/dx=150x-0.12x^2=0 x(150-0.12x)=0 x=0 or 150-0.12x=0 We have two x, we still don't know which one gives max y. Just plug it into the equation, and you'll get your max:D